What is the answer to this problem? I'm having a hard time with it?
A scalene triangle has three unequal sides. The perimeter of a scalene triangle is 144 m. If the first side is twice as long as the second side and the third
side is 24 m longer than the second side, then what are the measures of the sides?

Found 2 solutions by Earlsdon, dabanfield:
Answer by Earlsdon(6294)

Let the first side be A, the second side B, and the third side C.
1) A+B+C = 144 "The perimeter of a scalene triangle is 144 m".
2) A = 2B "...the first side (A) is twice as long as the second side (B)..."
3) C = B+24 "...the third side (C) is 24 m longer than the second side,..."
In equation 1) replace the A with equation 2) and the C with equation 3).
1a) (2B)+B+(B+24) = 144 Simplify.
1c) 4B+24 = 144 Subtract 24 from both sides.
1d) 4B = 120 Divide both sides by 4.
1e)B = 30
2a) A = 2B = 2(30) = 60
3a) C = B+24 = 30+24 = 54
The measures of the sides are:
A = 60 m.
B = 30 m.
C = 54 m.
Check:
A+B+C = 60+30+54 = 144 m (The given perimeter of the scalene triangle).

Answer by dabanfield(803)

A scalene triangle has three unequal sides. The perimeter of a scalene triangle is 144 m. If the first side is twice as long as the second side and the third
side is 24 m longer than the second side, then what are the measures of the sides?

Let x be the second side of the triangle above. Then the first side is 2*x and the third side is x + 24.
The perimeter is 144 = x + (2*x)+ (x+24)
Solve for x then substitute x for into 2*x and x+24 for the other sides.